OpenStudy (anonymous):

A.1+i and esquer 2(cos (pi/4)+i sin (pi/4)) B. 3+4 i and 5(cos (4/3)+i sin (4/3)) C.-3-3i,escuer2(cos (-pi/4)+i sin (-pi/4)) D. 1-i, -escuer2 (cos (-pi/4)+ i sin (-sin/4)) Which pair represents the same complex number?

OpenStudy (anonymous):

very hard to read. but \[i+i=\sqrt{2}(\cos(\frac{\pi}{4})+i\sin(\frac{\pi}{4}))\] so if that is a choice it is right

OpenStudy (anonymous):

i meant 1 + i of course